The table below gives the list price and the number of bids received for five randomly selected items sold through online aucStep 2 of 6: Find the estimated y-intercept. Round your answer to three decimal places. Answer(How to Enter) 2 Points E KeypaStep 3 of 6: Find the estimated value of y when x = 130. Round your answer to three decimal places. Answer(How to Enter) 2 PoStep 4 of 6: Substitute the values you found in steps 1 and 2 into the equation for the regression line to find the estimatedCopy Data Step 5 of 6: Find the error prediction when x = 130. Round your answer to three decimal places. Answer (How to EnteStep 6 of 6: Find the value of the coefficient of determination. Round your answer to three decimal places. Answer (How to EnSummation Table 1690 Bid 1 Bid 2 130 156 13 14 16900 24336 169 196 2184 Bid 3 168 16 2688 28224 256 Bid 4 178 18 3204 31684 3
Answer #1

The statistical software output for this problem is:

Simple linear regression results: Dependent Variable: Bids Independent Variable: Dollars Bids = -2.3832021 + 0.11264217 Dolla

Hence,

Step - 1: Slope = 0.113

Step - 2: y - Intercept = -2.383

Step - 3: Estimated value = 12.260

Step - 4: Change in dependent variable = b1 = 0.113

Step - 5: Error prediction = 13 - 12.260 = 0.740

Step - 6: Coefficient of determination = 0.892